# LATTE: Learning Aligned Transactions and Textual Embeddings for Bank Clients

### v1

@@ Abstract
Learning clients embeddings from sequences of their historic communications is central to financial applications. While large language models (LLMs) offer general world knowledge, their direct use on long event sequences is computationally expensive and impractical in real-world pipelines. In this paper, we propose LATTE, a contrastive learning framework that aligns raw event embeddings with semantic embeddings from frozen LLMs. Behavioral features are summarized into short prompts, embedded by the LLM, and used as supervision via contrastive loss. The proposed approach significantly reduces inference cost and input size compared to conventional processing of complete sequence by LLM. We experimentally show that our method outperforms state-of-the-art techniques for learning event sequence representations on real-world financial datasets while remaining deployable in latency-sensitive environments.
@@ 1 Introduction
Learning representations from structured event sequences [26, 27, 28, 29, 30] remains a core challenge in industrial applications of healthcare [36], education [12], e-commerce [10, 37], and, especially, finance [31, 35]. In the banking sector, hundreds of streams of temporally ordered, high-dimensional, and often sparse tabular data [41], such as transaction logs, payment histories, and customer interactions, are generated continuously. These data underpin a broad spectrum of business-critical tasks, including churn prediction, risk assessment, credit scoring, and personalized targeting [13]. Despite abundant customer interaction data, high-quality labels for such downstream tasks remain limited. This shortage of timely supervision hinders the scalability of supervised learning in production settings. It highlights the need for self-supervised approaches to derive robust and semantically rich representations directly from raw behavioral sequences [14]. Prevailing self-supervised approaches for modeling event sequences adopt such objectives as contrastive learning [1], next-event prediction [22], and latent sequence modeling techniques, e.g., Contrastive Predictive Coding (CPC) [23], aiming to capture temporal dependencies and user intent without relying on manual supervision.
@@
Figure 1: Figure of Merit comparing ROC-AUC performance and inference speed (samples/sec/GPU) on the Gender prediction task. Compared methods include LATTE [1], LATTE-S[1], CoLES [1] RTD [24], CPC [23], NSP [25], NPPR [22], DeTPP [33], IFTPP [34], IFTPP-T [34], ESQA [15], and TALLRec [19].
Recent advancements in Large Language Models (LLMs) offer new opportunities to enhance representation learning from event sequences. Trained on diverse and large-scale corpora, LLMs encode elements often implicit or absent in structured event datasets, e.g., rich semantic priors about behavioral patterns, temporal dynamics, and domain knowledge. Leveraging this external knowledge can significantly improve the quality of user representations, particularly in financial applications [43]. Motivated by this potential, recent studies have explored adaptations of LLMs to structured data. For example, TALLRec [19], LLM-TRSR [20], and HKFR [21] transfer rich text understanding abilities of LLMs to recommender systems; TabLLM [18] targets tabular classification tasks; TEST [17] and Time-LLM [16] address time series; while ESQA [15] applies LLMs to event-sequence question answering.
@@
However, a naive direct application of LLMs to serialized sequential tabular data incurs substantial computational overhead due to the large token counts per user. For example, typical banking transaction datasets often contain hundreds of records per user, each serialized into dozens of tokens, exceeding practical context window limits and increasing inference and training times. Figure 1 shows that LLM-based models such as TALLRec [19] and ESQA [15] do not exceed an inference speed of 10 users per second, which severely limits their applicability in banking production environments.
Existing methods to mitigate this issue fall into two main categories. The first reduces context length by summarizing user histories with general-purpose LLMs [21, 20], which risks losing domain-specific information critical for accurate modeling. The second class of methods attempts to bypass context length constraints by encoding sequences in non-textual formats [17, 16]. However, these representations often discard the semantic content present in item descriptions (e.g., transaction categories or merchant details). Moreover, as user histories grow longer, these models either incur increased computational cost or suffer performance degradation due to limited model capacity.
@@
To address these limitations, we propose LATTE  [^1] The source code is available at https://anonymous.4open.science/r/LATTE-3BD2/, a scalable framework for Learning Aligned Transactions and Textual Embeddings. Instead of feeding entire sequences into LLMs, we extract compact client-level statistics and use an instruction-tuned LLM to generate natural language summaries. These summaries serve as weak labels, aligned with pretrained sequence embeddings from a lightweight encoder via contrastive learning. At inference, LATTE supports two modes: a standalone encoder (LATTE-S) that retains LLM-level semantics without added overhead, and a combined encoder (LATTE) that fuses textual and structural representations.
@@
To evaluate trade-offs between performance and efficiency, we introduce a Figure of Merit (FOM) comparing ROC-AUC and inference speed on the standard banking dataset of gender prediction task [3] (Figure 1). We evaluate LATTE in two variants, combining two embedding strategies: structural-only (LATTE-S) and concatenated with textual features (LATTE). In terms of performance, all versions of LATTE outperform the typical baseline financial methods. The structural-only variant (LATTE-S) achieves a ROC-AUC of 0.899 on the Gender task with an inference speed of 162 samples/sec/GPU, surpassing LLM-based models such as TALLRec and ESQA, being over $14\times$ faster. Across all types of textual encoders, the combined variant (LATTE) consistently achieves the highest overall ROC-AUC.
@@ 2 Proposed Approach
We aim to improve the quality of representations learned from transactional event sequences by introducing an auxiliary modality derived from textual behavioral descriptions. To this end, we propose a three-stage pipeline, LATTE, illustrated in Figure 2, which maps raw transaction sequences into rich embeddings suitable for downstream tasks.
@@
Let $T_{i}={x_{1},x_{2},\dots,x_{n}}$ denote the transaction sequence for client $i$, where each $x_{j}$ contains timestamped categorical and numerical attributes (e.g., amount, merchant category). As shown in Fig. 2A, we first compute a vector of summary features $s_{i}$ that aggregates behavioral patterns over the sequence: frequency of activity, merchant diversity, transaction types, temporal coverage, and income-expense structure. Categorical statistics are transformed into meaningful textual descriptions rather than raw indices. See Appendix C for the prompt template and a sample generated description. This profile is then serialized into a textual prompt (Fig. 2B), which is passed to a frozen instruction-tuned LLM (Fig. 2C) to generate a natural language description $d_{i}$ of the client’s behavior (Fig. 2D). The prompt includes a system message and a structured representation of $s_{i}$ designed to elicit coherent and interpretable responses.
@@
Simultaneously, the raw sequence $T_{i}$ is processed by a GRU-based sequence encoder (Fig. 2G) trained under a self-supervised CoLES objective [1], where each training example consists of two overlapping subsequences (positives) and contrastive negatives from other clients. This produces the sequence embedding $z_{i}^{\text{seq}}$ optimized to capture client-specific behavioral dynamics. In parallel, the description $d_{i}$ is passed through a frozen multilingual sentence encoder (Fig. 2F), producing the text embedding $z_{i}^{\text{text}}$ (Fig. 2E) via mean pooling.
@@
The embeddings $z_{i}^{\text{seq}}$ and $z_{i}^{\text{text}}$ are then aligned using one of three cross-modal contrastive losses (Fig. 2I), while keeping the text encoder fixed. The updated transaction embeddings $\left(z_{i}^{\text{seq}}\right)^{{}^{\prime}}$ (Fig. 2H), aligned with the textual representations, are evaluated on downstream tasks such as churn prediction (Fig. 2J).
@@
To perform cross-modal alignment, we introduce three contrastive heads: Symmetric Softmax, Pairwise Binary, and Orthogonal Regularized. Each was inspired by prior work on multimodal learning [6, 7, 32]. These heads differ in their alignment geometry and regularization, and we refer to them as LATTE [1], LATTE [2], and LATTE [3], respectively. All three heads use frozen text encoders and update only the transaction encoder. Downstream tasks are evaluated using the resulting $\left(z_{i}^{\text{seq}}\right)^{{}^{\prime}}$, while the textual modality provides semantically grounded alignment by enriching the training signal with contextual knowledge of categorical attributes that the sequence encoder alone cannot interpret.
@@
LATTE [2]: Pairwise Binary Contrastive Head applies sigmoid-based supervision across all sequence-description pairs in the batch:
$\mathcal{L}_{\text{sigm}}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{N}\log\left(\frac{1}{1+e^{y_{ij}(-\tau z_{i}^{\text{seq}}\cdot z_{j}^{\text{text}}+b)}},\right)$ (2)
where $y_{ij}=1$ for $i=j$ (positive pairs), $y_{ij}=-1$ otherwise (negatives), and $b$ is a trainable bias.
@@ 3.1 Data
We evaluate our method on three real-world datasets containing anonymized credit card transaction sequences from major financial institutions. Each dataset comprises client-level sequences with numerical and categorical attributes (e.g., amount, merchant category, transaction type), and includes an unlabeled subset used exclusively for representation learning. Churn [2] includes approximately 10K Rosbank clients labeled by future inactivity. Gender [3] and Age Group [4], provided by Sberbank, contain 15K and 50K clients respectively, annotated with demographic labels.
@@ 3.2 Validation Strategy
Each dataset is split into disjoint training and test partitions by reserving 10% of the labeled clients for evaluation. The remaining 90% of labeled users, together with all available unlabeled users, are used for training the embedding models. To assess the quality of learned representations, we adopt a 5-fold cross-validation procedure. Specifically, the labeled portion of the training data is divided into five equal-sized folds. For each fold $v$, we: (1) train a LightGBM [38] classifier on embeddings from the remaining four folds, and (2) evaluate it on the held-out test fold, computing a downstream performance metric $M_{v}$. For binary classification tasks (churn, gender), we report ROC-AUC; for multiclass age prediction, we report classification accuracy. The final performance is summarized as $\mu_{M}\pm\sigma_{M}$, where $\mu_{M}$ is the mean and $\sigma_{M}$ is the standard deviation over all five folds.
@@ 3.3 Baselines
Event sequence models. CPC [23] learns to predict future representations from past context via a contrastive loss. CoLES [1] improves temporal consistency by aligning overlapping subsequences using InfoNCE. NPPR [22] employs autoregressive training with dual objectives: predicting the next and reconstructing the previous event from masked sequences. All models operate on raw event sequences without external modalities.
Temporal point process models. DeTPP [33] models event timing and types using parametric point processes. IFTPP and IFTPP-T [34] use Transformer and GRU backbones with combined MAE and classification losses.
Natural language processing objectives. RTD [24] randomly replaces 15% of event tokens and predicts the original token; NSP [25] extends BERT’s next sentence prediction to sequences of events.
LLM-based approaches. We adapt TALLRec [19] and HKFR [21] from recommender systems for user embedding extraction by fine-tuning LLMs on serialized user-item histories using next-token prediction. Embeddings are derived via mean pooling over the final layer, following recent best practices [46, 42]. Additionally, we replace older backbones (e.g., LLaMA 7B [49], ChatGLM-6B [48]) with LLaMA 3.2 3B [47] to leverage architectural improvements and efficiency.
@@ 4.1 Main Results
LATTE [1] 0.867 ± 0.004 0.666 ± 0.005 0.900 ± 0.005
LATTE [2] 0.868 ± 0.006 0.671 ± 0.005 0.901 ± 0.005
LATTE [3] 0.857 ± 0.003 0.678 ± 0.003 0.902 ± 0.006
Table 1: Performance of client embeddings on downstream tasks. Bold indicates best result; underline indicates second-best; double underline indicates third-best
Table 1 presents the performance of client embeddings on three downstream tasks: churn prediction (AUC), age group classification (accuracy), and gender prediction (AUC). While traditional self-supervised objectives such as CPC, RTD, and NSP lag behind stronger baselines like CoLES and TALLRec, several variants of our proposed LATTE framework demonstrate clear gains across tasks. LATTE [1] achieves the best result on churn prediction (0.868 AUC), while LATTE [3] leads on both age group classification (0.678 accuracy) and gender prediction (0.902 AUC). These improvements indicate that incorporating synthetic textual supervision leads to consistently stronger representations of transactional behavior.

### v4

@@ Abstract
Learning clients embeddings from sequences of their historic communications is central to financial applications. While large language models (LLMs) offer general world knowledge, their direct use on long event sequences is computationally expensive and impractical in real-world pipelines. In this paper, we propose LATTE, a contrastive learning framework that aligns raw event embeddings with description-based semantic embeddings from frozen LLMs. Behavioral features based on statistical user descriptions are summarized into short prompts, embedded by the LLM, and used as supervision via contrastive loss. The proposed approach significantly reduces inference cost and input size compared to the conventional processing of complete sequences by LLM. We experimentally show that our method outperforms state-of-the-art techniques for learning event sequence representations on real-world financial datasets while remaining deployable in latency-sensitive environments.
LATTE: Learning Aligned Transactions and Textual Embeddings for Bank Clients
Egor Fadeev1, Dzhambulat Mollaev1, Aleksei Shestov1, Omar Zoloev1, 3, Artem Sakhno1, Dmitry Korolev1, Ivan Kireev1, Andrey Savchenko1,2, Maksim Makarenko1 1Sber AI Lab, Moscow, Russia 2ISP RAS Research Center for Trusted Artificial Intelligence, Moscow, Russia 3NUST MISIS, Moscow, Russia
@@ 1 Introduction
Research in natural language processing (NLP) has traditionally focused on unstructured text Bagheri et al. (2023). In contrast, many industrial applications of healthcare Wang et al. (2024b), education Liu et al. (2023), e-commerce Dai et al. (2023); Liu et al. (2025), and, especially, finance Babaev et al. (2019); Luetto et al. (2025), generate hundreds of streams of structured event (temporally ordered, high-dimensional, and often sparse tabular) data Osin et al. (2025), such as transaction logs, payment histories, and customer interactions, which are sequential, high-dimensional, and sparse Zhang et al. (2023); Osin et al. (2025). These data underpin a broad spectrum of business-critical tasks, including churn prediction, risk assessment, credit scoring, and personalized targeting Mollaev et al. (2025). Transaction sequences differ from text in three key ways: they are much longer (thousands of events in open datasets, millions in proprietary banking logs) Mollaev et al. (2025), each event includes multiple categorical and numerical attributes Zhang et al. (2023), and the main tasks are classification or regression rather than broad semantic benchmarks Muennighoff et al. (2023).
@@
Figure 1: Figure of Merit comparing ROC-AUC performance and inference speed (samples/sec/GPU) on the Gender prediction task. Compared methods include LATTE [1], LATTE-S[1], CoLES Babaev et al. (2022) RTD Clark et al. (2019), CPC Oord et al. (2018), NSP Devlin et al. (2019), NPPR (Skalski et al., 2023), DeTPP Karpukhin and Savchenko (2024), IFTPP Shchur et al. (2020), IFTPP-T Shchur et al. (2020), ESQA Abdullaeva et al. (2024), and TALLRec Bao et al. (2023).
@@
Recent works on applying Large Language Models (LLMs) for structured data Shi et al. (2023); Yu et al. (2025) highlight that progress in this domain depends on methods adapted to the unique statistical and causal structure of event data, rather than direct transfer of techniques from NLP. Moreover, a direct application of LLMs to serialized sequential tabular data incurs substantial computational overhead due to the large token counts per user. For example, typical banking transaction datasets often contain hundreds of records per user, each serialized into dozens of tokens, exceeding practical context window limits and increasing inference and training times Shestov et al. (2025). Figure 1 shows that LLM-based models such as TALLRec Bao et al. (2023) and ESQA Abdullaeva et al. (2024) do not exceed an inference speed of 10 users per second, which severely limits their applicability in banking production environments.
@@
To address the limitations of existing techniques, we propose LATTE, a scalable framework for Learning Aligned Transactions and Textual Embeddings. Instead of feeding entire sequences into LLMs, we extract compact client-level statistics and use an instruction-tuned LLM to generate natural language summaries. These summaries serve as weak labels, aligned with pretrained sequence embeddings from a lightweight encoder via contrastive learning. At inference, LATTE supports two modes: a standalone encoder (LATTE-S) that retains LLM-level semantics without added overhead, and a combined encoder (LATTE) that fuses textual representations enriched with statistical features and structural embeddings.
@@
To evaluate trade-offs between performance and efficiency, we introduce a Figure of Merit (FOM) comparing ROC-AUC and inference speed on the standard banking dataset of gender prediction task (Sberbank, 2021a) (Figure 1). We evaluate LATTE in two variants, combining two embedding strategies: structural-only (LATTE-S) and concatenated with textual features (LATTE). In terms of performance, all versions of LATTE outperform the typical baseline financial methods. The structural-only variant (LATTE-S) achieves a ROC-AUC of 0.891 on the Gender task with an inference speed of 162 samples/sec/GPU, surpassing LLM-based models such as TALLRec and ESQA, being over $14\times$ faster. Across all types of textual encoders, the combined variant (LATTE) consistently achieves the highest overall ROC-AUC.
2 Related Works
Learning representations from structured event sequences Udovichenko et al. (2024); Yeshchenko and Mendling (2022); Kolosnjaji et al. (2016); Weiss and Hirsh (1998); Guo et al. (2020) remains a core challenge in industrial applications. Despite abundant customer interaction data, high-quality labels for event sequences in most typical downstream tasks (campaigning, churn prediction, etc.) remain limited Mollaev et al. (2025). This shortage of timely supervision hinders the scalability of supervised learning in production settings. It highlights the need for self-supervised approaches to derive robust and semantically rich representations directly from raw behavioral sequences Gui et al. (2024). Prevailing self-supervised approaches for modeling event sequences adopt such objectives as contrastive learning Babaev et al. (2022), next-event prediction Skalski et al. (2023), and latent sequence modeling techniques, e.g., Contrastive Predictive Coding (CPC) Oord et al. (2018), aiming to capture temporal dependencies and user intent without relying on manual supervision.
Appearance of LLMs offer new opportunities to enhance representation learning from event sequences. Trained on diverse and large-scale corpora, LLMs encode elements often implicit or absent in structured event datasets, e.g., rich semantic priors about behavioral patterns, temporal dynamics, and domain knowledge. Leveraging this external knowledge can significantly improve the quality of user representations, particularly in financial applications Ruan et al. (2024). Motivated by this potential, recent studies have explored adaptations of LLMs to structured data. For example, TALLRec Bao et al. (2023), LLM-TRSR Zheng et al. (2024), and HKFR Yin et al. (2023) transfer rich text understanding abilities of LLMs to recommender systems; TabLLM Hegselmann et al. (2023) targets tabular classification tasks; TEST Sun et al. (2024) and Time-LLM Jin et al. (2024) address time series; while ESQA Abdullaeva et al. (2024) applies LLMs to event-sequence question answering.
Existing methods to mitigate this issue fall into two main categories. The first reduces context length by summarizing user histories with general-purpose LLMs Yin et al. (2023); Zheng et al. (2024), which risks losing domain-specific information critical for accurate modeling. The second class of methods attempts to bypass context length constraints by encoding sequences in non-textual formats Sun et al. (2024); Jin et al. (2024). However, these representations often discard the semantic content present in item descriptions (e.g., transaction categories or merchant details). Moreover, as user histories grow longer, these models either incur increased computational cost or suffer performance degradation due to limited model capacity.
Recent advances in event sequence modeling reinforce the distinction between unstructured texts from NLP and structured data. Work on spatio-temporal clustering shows that standard neural point process models fail to capture hierarchical spatial structures and multi-type dependencies, requiring new architectures tailored to these properties Yu et al. (2025). Other studies demonstrate that even transformer-based approaches underperform when causal relations between event types are ignored, motivating causality-aware attention mechanisms Shou et al. (2023). Further results indicate that while Large Language Models (LLMs) can aid event prediction through abductive reasoning, they are effective only when combined with specialized sequence models Shi et al. (2023). Together, these findings highlight that progress in this domain depends on methods adapted to the unique statistical and causal structure of event data, rather than direct transfer of techniques from NLP.
@@ 3 Proposed Approach
We aim to improve the quality of representations learned from transactional event sequences by introducing an auxiliary textual modality that verbalizes statistical properties of user behavior. To this end, we propose a three-stage pipeline, LATTE, illustrated in Figure 2, which maps raw transaction sequences into rich embeddings suitable for downstream tasks.
@@
Let $T_{i}={x_{1},x_{2},\dots,x_{n}}$ denote the transaction sequence for client $i$, where each $x_{j}$ contains timestamped categorical and numerical attributes (e.g., amount, merchant category). As shown in Fig. 2a, we first compute a vector of summary features $s_{i}$ that aggregates behavioral patterns over the sequence: frequency of activity, merchant diversity, transaction types, temporal coverage, and income-expense structure. Behavioral features are then transformed into meaningful textual descriptions rather than raw indices, allowing these summaries to be further enriched with the semantic knowledge of the LLM. The prompt template and a sample generated description are presented in Appendix C. This profile is then serialized into a textual prompt (Fig. 2b), which is passed to a frozen instruction-tuned LLM (Fig. 2c) to generate a natural language description $d_{i}$ of the client’s behavior (Fig. 2d). The prompt includes a system message and a structured representation of $s_{i}$ designed to elicit coherent and interpretable responses.
@@
Simultaneously, the raw sequence $T_{i}$ is processed by a GRU-based sequence encoder (Fig. 2G) trained under a self-supervised CoLES objective (Babaev et al., 2022), where each training example consists of two overlapping subsequences (positives) and contrastive negatives from other clients. This produces the sequence embedding $z_{i}^{\text{seq}}$ optimized to capture client-specific behavioral dynamics. In parallel, the description $d_{i}$ is passed through a frozen multilingual sentence encoder (Fig. 2f) with mean pooling, producing the text embedding $z_{i}^{\text{text}}$ enriched with raw statistical features derived from the original sequence.
@@
The embeddings $z_{i}^{\text{seq}}$ and $z_{i}^{\text{text}}$ are then aligned using one of three cross-modal contrastive losses (Fig. 2i), while keeping the text encoder fixed. The updated transaction embeddings $\left(z_{i}^{\text{seq}}\right)^{{}^{\prime}}$ (Fig. 2h), aligned with the textual representations, are evaluated on downstream tasks such as churn prediction (Fig. 2j). Appendix A provides additional details regarding the pipeline.
@@
To perform cross-modal alignment, we introduce two different contrastive heads: Symmetric Softmax and Orthogonal Regularized. Each was inspired by prior work on multimodal learning (Radford et al., 2021; Jiang et al., 2023). These heads differ in their alignment geometry and regularization, and we refer to them as LATTE [1], and LATTE [2], respectively. Both heads use frozen text encoders and update only the transaction encoder. Downstream tasks are evaluated using the resulting $\left(z_{i}^{\text{seq}}\right)^{{}^{\prime}}$, while the textual modality provides semantically grounded alignment by enriching the training signal with contextual knowledge of categorical attributes that the sequence encoder alone cannot interpret.
@@ 4.1 Data
We evaluate our method on three real-world datasets containing anonymized credit card transaction sequences from major financial institutions. Each dataset comprises client-level sequences with numerical and categorical attributes (e.g., amount, merchant category, transaction type), and includes an unlabeled subset used exclusively for representation learning. Churn (Rosbank, 2021) includes approximately 10K Rosbank clients labeled by future inactivity. Gender (Sberbank, 2021a) and Age Group (Sberbank, 2021b), provided by Sberbank, contain 15K and 50K clients respectively, annotated with demographic labels.
@@ 4.2 Validation Strategy
Each dataset is split into disjoint training and test partitions by reserving 10% of the labeled clients for evaluation. The remaining 90% of labeled users, together with all available unlabeled users, are used for training the embedding models. To assess the quality of learned representations, we adopt a 5-fold cross-validation procedure. Specifically, the labeled portion of the training data is divided into five equal-sized folds. For each fold $v$, we: (1) train a LightGBM (Ke et al., 2017) classifier on embeddings from the remaining four folds, and (2) evaluate it on the held-out test fold, computing a downstream performance metric $M_{v}$. For binary classification tasks (churn, gender), we report ROC-AUC; for multiclass age prediction, we report classification accuracy. The final performance is summarized as $\mu_{M}\pm\sigma_{M}$, where $\mu_{M}$ is the mean and $\sigma_{M}$ is the standard deviation over all five folds.
@@ 4.3 Baselines
Event sequence models. CPC (Oord et al., 2018) learns to predict future representations from past context via a contrastive loss. CoLES (Babaev et al., 2022) improves temporal consistency by aligning overlapping subsequences using InfoNCE. NPPR (Skalski et al., 2023) employs autoregressive training with dual objectives: predicting the next and reconstructing the previous event from masked sequences. All models operate on raw event sequences without external modalities.
Temporal point process models. DeTPP (Karpukhin and Savchenko, 2024) models event timing and types using parametric point processes. IFTPP and IFTPP-T (Shchur et al., 2020) use Transformer and GRU backbones with combined MAE and classification losses.
Natural language processing objectives. RTD (Clark et al., 2019) randomly replaces 15% of event tokens and predicts the original token; NSP (Devlin et al., 2019) extends BERT’s next sentence prediction to sequences of events.
LLM-based approaches. We adapt TALLRec Bao et al. (2023) and HKFR Yin et al. (2023) from recommender systems for user embedding extraction by fine-tuning LLMs on serialized user-item histories using next-token prediction. Embeddings are derived via mean pooling over the final layer, following recent best practices BehnamGhader et al. (2024); Muennighoff et al. (2024). Additionally, we replace older backbones (e.g., LLaMA 7B Touvron et al. (2023), ChatGLM-6B Du et al. (2022)) with LLaMA 3.2 3B Touvron et al. (2024) to leverage architectural improvements and efficiency.
@@ 5.1 Main Results
LATTE [1] 0.869 ± 0.004 0.665 ± 0.005 0.900 ± 0.005
LATTE [2] 0.872 ± 0.004 0.663 ± 0.003 0.898 ± 0.006
Table 1: Performance of client embeddings on downstream tasks. Bold indicates best result; underline indicates second-best
Table 1 presents the performance of client embeddings on three downstream tasks: churn prediction (AUC), age group classification (accuracy), and gender prediction (AUC). While traditional self-supervised objectives such as CPC, RTD, and NSP lag behind stronger baselines like CoLES and TALLRec, several variants of our proposed LATTE framework demonstrate clear gains across tasks. LATTE [2] achieves the best result on churn prediction (0.872 AUC), while LATTE [1] leads on both age group classification (0.665 accuracy) and gender prediction (0.900 AUC). These improvements indicate that incorporating statistic-based textual supervision leads to consistently stronger representations of transactional behavior.